Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Letters (probability)

Expert replies
by LSB » Mon Sep 01, 2008 11:33 am
Q from GMAT Prep.

Is this not a dependent prob Q? Please post your response w explanation. Appreciate it guys
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by LSB » Mon Sep 01, 2008 1:01 pm
the attachments in this post are quite large. I am reattaching as word documents.
You do not have the required permissions to view the files attached to this post.
Join the discussion

by parallel_chase » Mon Sep 01, 2008 1:17 pm
The probability of A letter going to A envelop and B, C, and D in wrong

A going into right envelop is 1/4
B going into wrong envelop is 2/3 ( 1-1/3)
C going into wrong envelop is 1/2
D automatically will go into the wrong envelop 1

1/4*2/3*1/2*1 = 1/12 (multiplication because everything occurs at once)

Similarly the same procedure will follow for B, C & D and the result will also be the same

1/12+1/12+1/12+1/12 = 4/12 = 1/3.
Join the discussion

by 4meonly » Wed Sep 03, 2008 7:25 am
Anybody, please find the flaw in my reasosning:

1. To make 1st letter with correct adress:
1/4*2/3*1/2*1 = 1/12

2. To make 2nd letter with correct adress:
3/4*1/3*1/2*1 = 1/8

3. To make 3rd letter with correct adress:
3/4*2/3*1/2*1 = 1/4

4. To make 4th letter with correct adress:
3/4*2/3*1/2*1 = 1/4

summ will be 17/24
Join the discussion

by canuckclint » Wed Sep 03, 2008 4:15 pm
4meonly wrote:Anybody, please find the flaw in my reasosning:

1. To make 1st letter with correct adress:
1/4*2/3*1/2*1 = 1/12

2. To make 2nd letter with correct adress:
3/4*1/3*1/2*1 = 1/8

3. To make 3rd letter with correct adress:
3/4*2/3*1/2*1 = 1/4

4. To make 4th letter with correct adress:
3/4*2/3*1/2*1 = 1/4

summ will be 17/24
The flaw is in 1 and multiplications in general
1 should be = 2/24
Join the discussion

by 4meonly » Thu Sep 04, 2008 8:00 am
1 should be = 2/24

does not this mean that 2/24 = 1/12 ?
Join the discussion